# Learning with Calibration: Exploring Test-Time Computing of Spatio-Temporal Forecasting

### v1

@@ 1 Introduction
Spatio-temporal forecasting (STF) aims to predict the future state of dynamic systems from historical spatio-temporal observations and underpins many real-world applications, such as traffic flow forecasting guo2021learning , air quality forecasting nguyen2023climax , and energy consumption forecasting wang2019review . Although spatio-temporal neural networks shao2024exploring ; jin2023spatio ; jin2024survey , which couple spatial neural operators with temporal neural operators, have achieved remarkable progress on these tasks, their deployment in practical environments remains fraught with challenges. These observations, typically collected by sensors, are frequently corrupted by noise, outliers (e.g., spikes or dropouts due to hardware failure) xu2024spatiotemporal , and more commonly, non-stationary distribution shifts arising from sensor aging and seasonal patterns wang2024evaluating .
To enhance generalization and performance, prior work has focused primarily on out-of-distribution (OOD) learning for ST data during the training phase: designing architectures that resist perturbations ji2025seeing ; ma2024stop ; wang2024stone ; zhou2024coms2t ; zhou2023maintaining ; xia2023deciphering , augmenting training data with noise or adversarial examples bai2025alleviating ; zhang2023automated ; liu2022practical , and introducing specialized loss functions or regularizers li2024flashst ; zhang2023spatial to counteract distribution drift. However, these methods share fundamental limitations: they assume that the training data sufficiently captures all future target domain invariance—a premise that is rarely valid in real-world settings. Concurrently, an emerging paradigm of continual fine-tuning chen2025eac ; yi2024get ; wang2024towards ; lee2024continual ; wang2023pattern ; wang2023knowledge ; kieu2024team ; chen2021trafficstream has become popular in spatio-temporal learning by continuously tuning the model to adapt to dynamic changes. Though promising, it still divides the target domain into multiple periods of training and testing and relies on period-specific training data to optimize model, thereby failing in data-scarse scenarios.
@@
Source Target Train-Time Test-Time
OOD Learning STONE wang2024stone , CaST xia2023deciphering $\langle\mathbf{X}^{s},\mathbf{Y}^{s}\rangle$ $\min_{f_{\theta}}\;\max_{e^{*}\in\mathcal{E}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{s},\mathbf{Y}^{s}\mid e^{*})}\bigl{[}L(f_{\theta}(x),y)\bigr{]}$
Continual Fine-Tuning EAC chen2025eac , TrafficStream chen2021trafficstream $\langle\mathbf{X}^{t},\mathbf{Y}^{t}\rangle$ $\min_{f_{{\theta}^{\tau}}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t},\mathbf{Y}^{t}\mid e^{\tau})}\bigl{[}L(f_{{\theta}^{\tau}}(x),y)\bigr{]}$
Test-Time Training TTT-ST chen2024test $\mathbf{X}^{s}$ $\mathbf{X}^{t}$
@@
Online Continual Learning DOST wang2024distribution $\mathbf{X}^{t}$ $\min_{f_{\theta(\delta)}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl{[}L(f_{\theta(\delta)}(x),y)\bigr{]}$
Test-Time Computing ST-TTC (Our) $\mathbf{X}^{t}$ $\min_{g_{\theta}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl{[}L\bigl{(}g_{\theta}(f_{\theta}(x)),y\bigr{)}\bigr{]}$
@@
Recently, leveraging test-time information has attracted widespread attention for its ability to significantly improve language model performance on complex reasoning tasks snell2025scaling ; akyurek2024surprising . In computer vision, this concept has already been extensively developed: test-time training (TTT) was first introduced by sun2020test , which defines an auxiliary self-supervised task applied to both training and test samples to better balance bias and variance gandelsman2022test . A similar idea was adapted to spatio-temporal forecasting in TTT-ST chen2024test . Unlike language and vision settings—where obtaining ground-truth labels for test samples at inference time is nearly impossible—STF benefits from label autocorrelation cressie2011statistics : each observation strongly depends on its predecessor, and training instances are constructed from sliding windows, which provide access to historical samples and their true labels. Moreover, this property makes STF also require timeliness shao2024stemo , that is, the additional computing time during inference must be less than the window-stride interval. A recent STF method, DOST wang2024distribution , explores online continual learning, which initially explored this direction. It uses historical test sample labels to dynamically adapt the modified model architecture. Though promising, these approaches typically involve complex self-supervised tasks or structural adaptations and still fall short of the timeliness demands of STF.
@@ 2 Related Work
Spatio-Temporal Forecasting. Spatio-temporal sequences can be regarded as spatially extended multivariate time series. Although one can trivially apply multivariate forecasting methods nie2023time ; zeng2023transformers ; biller2003modeling ; box1970distribution independently at each location, such decoupling of spatial and temporal dependencies invariably yields suboptimal results shao2024exploring . Classical spatio-temporal forecasting method instead relies on shallow models or spatio-temporal kernels, including feature-based methods ohashi2012wind ; zheng2015forecasting , state space models cliff1975space ; pfeifer1980starima ; bahadori2014fast , and Gaussian process models flaxman2015machine ; senanayake2016predicting . Unfortunately, the overall nonlinearity of these models is limited, and the high complexity of computation and storage further hinders the availability of massive training instances shi2018machine . In recent years, spatio-temporal neural networks kumar2024spatio ; jin2024survey ; jin2023spatio have been widely adopted to learn the complex dynamics of such systems. Early work concentrated on devising neural operators to extract spatial or temporal correlation yu2018spatio ; wu2019graph ; shao2022spatial ; liu2023spatio ; fang2025efficient and on designing fusion architectures to integrate them li2018diffusion ; guo2019attention ; oreshkin2021fc ; cini2023taming ; deng2024disentangling . More recent efforts have explored domain-invariant representation learning ma2024stop ; wang2024stone ; xia2023deciphering and continual model adaptation chen2025eac ; wang2023pattern ; chen2021trafficstream to better accommodate unseen environmental shifts. However, these methods still depend exclusively on offline training data and thus cannot deliver truly timely and effective adaptation in real settings.
@@
Test-Time Computing. Test-time computation is inspired by the human cognition kahneman2011thinking , in which additional computational effort is allocated during inference to improve task performance. This insight has recently driven considerable interest in the nature language process community, fueled by the success of reasoning-augmented language models (e.g., o1 jaech2024openai and r1 guo2025deepseek ) that activate and adapt internal computations at test time via supervised fine-tuning or reinforcement learning (RL) zhang2025and . While the generalization properties of RL-based adaptation remain debated yue2025does , the notion of supervised learning on unlabeled test data dates back to “transductive learning” Gammerman98lbyt in the 1990s and has demonstrated empirical benefits vapnik1999nature ; bottou1992local . In the computer vision domain, this idea was formalized as Test-Time Training sun2020test , which attaches an auxiliary self-supervised head to enable online adaptation to each test instance—a paradigm subsequently generalized as test-time adaptation wangtent ; jangtest ; wang2022continual ; liang2025comprehensive . However, spatio-temporal forecasting has seen limited exploration of such techniques. TTT-ST chen2024test applies TTT-style auxiliary objectives during training and continues to update at inference, and DOST wang2024distribution further incorporates dynamic learning mechanisms within modified model architectures for test-time updates. In addition, some methods zhang2023test ; guo2024online are conceptually close to ours, such as CompFormer zhang2023test , which proposes a test-time compensated representation learning framework, but still requires access to additional training data. Notably, we formalize the test-time computing of spatio-temporal forecasting, and propose a unified learning-with-calibration framework that is general, lightweight, efficient, and effective for STF at test-time.
@@ 3 Preliminaries
Problem Definition. Let $x\in\mathbb{R}^{T\times C}$ denote the multivariate time series recorded at each location sensor, capturing the dynamic observations of $C$ measured features in $T$ consecutive time steps. Stacking these sequences for all $N$ locations yields the spatio-temporal tensor $X\in\mathbb{R}^{N\times T\times C}$. Given historical observations $X^{h}\in\mathbb{R}^{N\times T^{h}\times C}$ (and an optional spatial correlation graph $\mathcal{G}$ representing the spatial relationships of $N$ locations), spatio-temporal forecasting aims to learn a mapping $f_{\theta}:(X^{h},\mathcal{G})\longmapsto X^{f}\in\mathbb{R}^{N\times T^{f}\times C}$, where $X^{f}$ is the signal for the next $T^{f}$ time steps. In practice, according to shao2024exploring ; li2018diffusion , the feature to be predicted is usually only the target variable.
@@ 4 Methodology
4.1 What to Compute? Spectral Domain Calibrator with Phase-Amplitude Modulation
Motivation. Spatio-temporal data, such as traffic flow and air quality, often exhibit periodic patterns (e.g., daily or weekly cycles). However, in real-world deployments, these patterns are not stationary; they are dynamically influenced by various internal and external factors wang2024robust . Such influences lead to non-stationarities manifesting as fluctuations in amplitude (e.g., increased or decreased traffic peaks due to seasonal changes) or phase shifts (e.g., peak hours are advanced or delayed due to traffic congestion). Pre-training models typically fit fixed periodic patterns during training, which makes them vulnerable to performance degradation under such persistent dynamic changes during inference wang2024evaluating . Therefore, we argue that the goal of test-time computation is: how to design an effective calibrator that can efficiently capture such gradual systematic bias from the pattern to correct the prediction errors caused by non-stationarity, while avoiding overfitting to random noise?
Key Challenges. While correction in the time domain is possible zhang2023test ; guo2024online , it often requires extensive parameterization, leading to increased model complexity and limited ability to capture evolving periodic structures. Moreover, the coupled structural and branching modules wang2024distribution ; chen2024test are prone to overfitting the random noise in the spatio-temporal evolution. To address this, we propose calibration in the spectral domain, where periodic variations are more transparently expressed as changes in the amplitude and phase of specific frequency components. Spectral correction offers a potentially more direct and robust solution. However, this introduces two main challenges: ❶ the degree of non-stationarity varies across spatial nodes; and ❷ full-spectrum parameterization is computationally expensive. The core problem thus becomes how to design a lightweight, spatial-aware calibrator.
@@ 4.1 What to Compute? Spectral Domain Calibrator with Phase-Amplitude Modulation
  • Group-wise Modulation. To ensure lightweight and balanced spectrum expression, we divide the $M$ frequency bins into $G$ contiguous groups of size $\lfloor M/G\rfloor$, and learn per-group, per-node amplitude and phase offsets $\lambda^{\alpha}\in\mathbb{R}^{G\times N\times 1},\lambda^{\phi}\in\mathbb{R}^{G\times N\times 1}$. For each group $g\in\{1,\dots,G\}$, we apply $A_{g}^{{}^{\prime}}=A_{g}\odot(1+\lambda^{\alpha}_{g}\bigr{)},P_{g}^{{}^{\prime}}=P_{g}+\lambda^{\phi}_{g},$ and reconstruct the spectrum as $Y_{f}^{{}^{\prime}}=\bigcup_{g=1}^{G}A^{{}^{\prime}}_{g}\odot e^{(j\,P^{{}^{\prime}}_{g})}$.
  • Inverse Transform. Finally, the calibrated time-domain signal is obtained by Inverser rFFT $\hat{y}_{cal}=\mathrm{irFFT}\bigl{(}Y_{f})\;\in\;\mathbb{R}^{B\times N\times T},$ along the frequency dimension.
For clarity, we provide a Algorithm workflow 1 and Pytorch-Style Pseudocode 2 in Appendix C.1.
Complexity Analysis. The full-spectrum parameterization learns independent amplitude and phase offsets for each of the $M=T/2+1$ frequency bins and $N$ nodes, totaling $2NM$ parameters. In contrast, our $G$-group design learns only $2NG$ parameters. Since $G$ is a constant and $M$ grows linearly with $T$, $G\ll M$ is usually the case. For large-scale long-term scenario, this significantly reduces memory footprint and gradient update cost while retaining interpretable per-band calibration.
Theoretical Analysis. We also provide a theoretical bound on the output perturbation induced by the SD-Calibrator, ensuring controlled deviation from the original prediction to prevent overfitting (Please refer to Theorem 1 and the proof in Appendix B).
4.2 How to Compute? Flash Gradient Update with Streaming Memory Queue

### v2

@@ 1 Introduction
Spatio-temporal forecasting (STF) aims to predict the future state of dynamic systems from historical spatio-temporal observations and underpins many real-world applications, such as traffic flow forecasting (Guo et al., 2021), air quality forecasting (Nguyen et al., 2023), and energy consumption forecasting (Wang et al., 2019). Although spatio-temporal neural networks (Jin et al., 2023, 2024; Shao et al., 2024b), which couple spatial neural operators with temporal neural operators, have achieved remarkable progress on these tasks, their deployment in practical environments remains fraught with challenges. These observations, typically collected by sensors, are frequently corrupted by noise, outliers (e.g., spikes or dropouts due to hardware failure) (Xu et al., 2024a), and more commonly, non-stationary distribution shifts arising from sensor aging and seasonal patterns (Wang et al., 2024d).
To enhance generalization and performance, prior work has focused primarily on out-of-distribution (OOD) learning for ST data during the training phase: designing architectures that resist perturbations (Ji et al., 2025; Ma et al., 2024; Wang et al., 2024a; Xia et al., 2023), augmenting training data with noise or adversarial examples (Bai et al., 2025; Liu et al., 2022; Zhang et al., 2023a), and introducing specialized loss functions or regularizers (Li et al., 2024b; Zhang et al., 2023b) to counteract distribution drift. However, these methods share fundamental limitations: they assume that the training data sufficiently captures all future target domain invariance, a premise that is rarely valid in real-world settings. Concurrently, an emerging paradigm of continual fine-tuning (Chen and Liang, 2025; Chen et al., 2021; Lee and Park, 2024; Wang et al., 2024b, 2023a, 2023b) has become popular in spatio-temporal learning by continuously tuning the model to adapt to dynamic changes. Though promising, it still divides the target domain into multiple periods of training and testing and relies on period-specific training data to optimize model, thereby failing in data-scarse scenarios.
@@
Source Target Train‐Time Test‐Time
OOD Learning STONE (Wang et al., 2024a), CaST (Xia et al., 2023) $\langle\mathbf{X}^{s},\mathbf{Y}^{s}\rangle$ $\min_{f_{\theta}}\;\max_{e^{*}\in\mathcal{E}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{s},\mathbf{Y}^{s}\mid e^{*})}\bigl[L(f_{\theta}(x),y)\bigr]$
Continual Fine‐Tuning EAC (Chen and Liang, 2025), TrafficStream (Chen et al., 2021) $\langle\mathbf{X}^{t},\mathbf{Y}^{t}\rangle$ $\min_{f_{{\theta}^{\tau}}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t},\mathbf{Y}^{t}\mid e^{\tau})}\bigl[L(f_{{\theta}^{\tau}}(x),y)\bigr]$
Test‐Time Training TTT-ST (Chen et al., 2024) $\mathbf{X}^{s}$ $\mathbf{X}^{t}$
@@
Online Continual Learning DOST (Wang et al., 2024c) $\mathbf{X}^{t}$ $\min_{f_{\theta(\delta)}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl[L(f_{\theta(\delta)}(x),y)\bigr]$
Test‐Time Computing ST-TTC (Our) $\mathbf{X}^{t}$ $\min_{g_{\theta}}\;\mathbb{E}_{(x,y)\sim P(\mathbf{X}^{t}\mid e)}\bigl[L\bigl(g_{\theta}(f_{\theta}(x)),y\bigr)\bigr]$
@@
Recently, leveraging test-time information has attracted widespread attention for its ability to significantly improve language model performance on complex reasoning tasks (Akyürek et al., 2024; Snell et al., 2025). In computer vision, this concept has already been extensively developed: test-time training (TTT) was first introduced by (Sun et al., 2020), which defines an auxiliary self-supervised task applied to both training and test samples to better balance bias and variance (Gandelsman et al., 2022). A similar idea was adapted to spatio-temporal forecasting in TTT-ST (Chen et al., 2024). Unlike language and vision settings—where obtaining ground-truth labels for test samples at inference time is nearly impossible—STF benefits from label autocorrelation (Cressie and Wikle, 2011): each observation strongly depends on its predecessor, and training instances are constructed from sliding windows, which provide access to historical samples and their true labels. Moreover, this property makes STF also require timeliness (Shao et al., 2024a), that is, the additional computing time during inference must be less than the window-stride interval. A recent STF method, DOST (Wang et al., 2024c), explores online continual learning, which initially explored this direction. It uses historical test sample labels to dynamically adapt the modified model architecture. Though promising, these approaches typically involve complex self-supervised tasks or structural adaptations and still fall short of the timeliness demands of STF.
@@ 2 Related Work
Spatio-Temporal Forecasting. Spatio-temporal sequences can be regarded as spatially extended multivariate time series. Although one can trivially apply multivariate forecasting methods (Biller and Nelson, 2003; Box and Pierce, 1970; Nie et al., 2023; Zeng et al., 2023) independently at each location, such decoupling of spatial and temporal dependencies invariably yields suboptimal results (Shao et al., 2024b). Classical spatio-temporal forecasting method instead relies on shallow models or spatio-temporal kernels, including feature-based methods (Ohashi and Torgo, 2012; Zheng et al., 2015), state space models (Bahadori et al., 2014; Cliff and Ord, 1975; Pfeifer and Deutsch, 1980), and Gaussian process models (Flaxman, 2015; Senanayake et al., 2016). Unfortunately, the overall nonlinearity of these models is limited, and the high complexity of computation and storage further hinders the availability of massive training instances (Shi and Yeung, 2018). In recent years, spatio-temporal neural networks (Jin et al., 2023, 2024; Kumar et al., 2024) have been widely adopted to learn the complex dynamics of such systems. Early work concentrated on devising neural operators to extract spatial or temporal correlation (Fang et al., 2025; Liu et al., 2023a; Shao et al., 2022b; Wu et al., 2019; Yu et al., 2018) and on designing fusion architectures to integrate them (Cini et al., 2023; Deng et al., 2024; Guo et al., 2019; Li et al., 2018; Oreshkin et al., 2021). More recent efforts have explored domain-invariant representation learning (Ma et al., 2024; Zhou et al., 2024, 2023) and continual model adaptation (Chen and Liang, 2025; Kieu et al., 2024; Chen et al., 2021; Yi et al., 2024) to better accommodate unseen environmental shifts. However, these methods still depend exclusively on offline training data and thus cannot deliver truly timely and effective adaptation in real settings.
@@
Test-Time Computing. Test-time computation is inspired by the human cognition (Kahneman, 2011), in which additional computational effort is allocated during inference to improve task performance. This insight has recently driven considerable interest in the nature language process community, fueled by the success of reasoning-augmented language models (e.g., o1 (Jaech et al., 2024) and r1 (Guo et al., 2025)) that activate and adapt internal computations at test time via supervised fine-tuning or reinforcement learning (RL) (Zhang et al., 2025). While the generalization properties of RL-based adaptation remain debated (Yue et al., 2025), the notion of supervised learning on unlabeled test data dates back to “transductive learning” (Gammerman et al., 1998) in the 1990s and has demonstrated empirical benefits (Bottou and Vapnik, 1992; Vapnik, 1999). In the computer vision domain, this idea was formalized as Test-Time Training (Sun et al., 2020), which attaches an auxiliary self-supervised head to enable online adaptation to each test instance—a paradigm subsequently generalized as test-time adaptation (Jang et al., 2023; Liang et al., 2025; Wang et al., 2021, 2022). However, spatio-temporal forecasting has seen limited exploration of such techniques. TTT-ST (Chen et al., 2024) applies TTT-style auxiliary objectives during training and continues to update at inference, and DOST (Wang et al., 2024c) further incorporates dynamic learning mechanisms within modified model architectures for test-time updates. In addition, some methods (Guo et al., 2024; Zhang et al., 2023c) are conceptually close to ours, such as CompFormer (Zhang et al., 2023c), which proposes a test-time compensated representation learning framework, but still requires access to additional training data. Notably, we formalize the test-time computing of spatio-temporal forecasting, and propose a unified learning-with-calibration framework that is general, lightweight, efficient, and effective for STF at test-time.
@@ 3 Preliminaries
Problem Definition. Let $x\in\mathbb{R}^{T\times C}$ denote the multivariate time series recorded at each location sensor, capturing the dynamic observations of $C$ measured features in $T$ consecutive time steps. Stacking these sequences for all $N$ locations yields the spatio-temporal tensor $X\in\mathbb{R}^{N\times T\times C}$. Given historical observations $X^{h}\in\mathbb{R}^{N\times T^{h}\times C}$ (and an optional spatial correlation graph $\mathcal{G}$ representing the spatial relationships of $N$ locations), spatio-temporal forecasting aims to learn a mapping $f_{\theta}:(X^{h},\mathcal{G})\longmapsto X^{f}\in\mathbb{R}^{N\times T^{f}\times C}$, where $X^{f}$ is the signal for the next $T^{f}$ time steps. In practice, according to (Li et al., 2018; Shao et al., 2024b), the feature to be predicted is usually only the target variable.
@@ 4 Methodology
4.1 What to Compute ?? Spectral Domain Calibrator with Phase-Amplitude Modulation
Motivation. Spatio-temporal data, such as traffic flow and air quality, often exhibit periodic patterns (e.g., daily or weekly cycles). However, in real-world deployments, these patterns are not stationary; they are dynamically influenced by various internal and external factors (Wang et al., 2024e). Such influences lead to non-stationarities manifesting as fluctuations in amplitude (e.g., increased or decreased traffic peaks due to seasonal changes) or phase shifts (e.g., peak hours are advanced or delayed due to traffic congestion). Pre-training models typically fit fixed periodic patterns during training, which makes them vulnerable to performance degradation under such persistent dynamic changes during inference (Wang et al., 2024d). Therefore, we argue that the goal of test-time computation is: how to design an effective calibrator that can efficiently capture such gradual systematic bias from the pattern to correct the prediction errors caused by non-stationarity, while avoiding overfitting to random noise?
Key Challenges. While correction in the time domain is possible (Guo et al., 2024; Zhang et al., 2023c), it often requires extensive parameterization, leading to increased model complexity and limited ability to capture evolving periodic structures. Moreover, the coupled structural and branching modules (Chen et al., 2024; Wang et al., 2024c) are prone to overfitting the random noise in the spatio-temporal evolution. To address this, we propose calibration in the spectral domain, where periodic variations are more transparently expressed as changes in the amplitude and phase of specific frequency components. Spectral correction offers a potentially more direct and robust solution. However, this introduces two main challenges: ❶ the degree of non-stationarity varies across spatial nodes; and ❷ full-spectrum parameterization is computationally expensive. The core problem thus becomes how to design a lightweight, spatial-aware calibrator.
@@ 4.1 What to Compute ?? Spectral Domain Calibrator with Phase-Amplitude Modulation
  • Group-wise Modulation. To ensure lightweight and balanced spectrum expression, we divide the $M$ frequency bins into $G$ contiguous groups of size $\lfloor M/G\rfloor$, and learn per‐group, per‐node amplitude and phase offsets $\lambda^{\alpha}\in\mathbb{R}^{G\times N\times 1},\lambda^{\phi}\in\mathbb{R}^{G\times N\times 1}$ (Note: Both $\lambda^{\alpha}$ and $\lambda^{\phi}$ are initialized to 0 to avoid incorrect calibration of predictions before learning). For each group $g\in\{1,\dots,G\}$, we apply $A_{g}^{{}^{\prime}}=A_{g}\odot(1+\lambda^{\alpha}_{g}\bigr),P_{g}^{{}^{\prime}}=P_{g}+\lambda^{\phi}_{g},$ and reconstruct the spectrum as $Y_{f}^{{}^{\prime}}=\bigcup_{g=1}^{G}A^{{}^{\prime}}_{g}\odot e^{(j\,P^{{}^{\prime}}_{g})}$.
  • Inverse Transform. Finally, the calibrated time‐domain signal is obtained by Inverser rFFT $\hat{y}_{cal}=\mathrm{irFFT}\bigl(Y_{f})\;\in\;\mathbb{R}^{B\times N\times T},$ along the frequency dimension.
For clarity, we provide a Algorithm workflow 1 and Pytorch-Style Pseudocode LABEL:pytorch_sdc in Appendix C.1.
Complexity Analysis. The full‐spectrum parameterization learns independent amplitude and phase offsets for each of the $M=T/2+1$ frequency bins and $N$ nodes, totaling $2NM$ parameters. In contrast, our GG‐group design learns only $2NG$ parameters. Since $G$ is a constant and $M$ grows linearly with $T$, $G\ll M$ is usually the case. For large-scale long-term scenario, this significantly reduces memory footprint and gradient update cost while retaining interpretable per-band calibration.
Theoretical Analysis. We also provide a theoretical approximate bound on the output perturbation induced by the SD-Calibrator, ensuring controlled deviation from the original prediction to prevent overfitting (Please refer to Theorem 1 and the proof in Appendix B).
4.2 How to Compute ?? Flash Gradient Update with Streaming Memory Queue